The Universal Operating System

One of the longest-established Linux distributions, Debian GNU/Linux is developed cooperatively by more than 800 volunteers located worldwide. Debian GNU/Linux is 100% Free Software, and will remain so—our commitment to freedom is demonstrated by the Debian Free Software Guidelines, now better known as the Open Source Definition.

Debian has a reputation for security and reliability—all known bugs in Debian are published in our bug tracking system, and we feed our changes back to the rest of the Linux community. “Big name” companies have recognised our technical excellence: Corel and Stormix based their Linux distributions on Debian, while VA research pays some of our developers to improve Debian as their full-time job.

Debian is a world-class operating system designed for those who want the best and aren’t afraid of learning how to use it:

• If you hadn’t spent time learning to drive a car, you’d have to walk or take the bus

• Would you rather use your computer at walking pace, or learn to drive an operating system that lets you get your work done more efficiently?

Debian GNU/Linux 2.2 (codenamed “potato”) is available now for Intel-based PCs and compatibles (“i386”), 68030-based Amigas, Atari STs, and Macs (“m68k”), Netwinders, Archimedes and others based on the ARM processor, newer Macs based on the PowerPC processor, and PCs and workstations based on Alpha and Sparc processors. Distributions for the HPPA, IA64, Mips, IBM S/390 and SH processors are under development. We are also working on our long-term project Debian GNU/HURD, the first ever operating system distribution based on the GNU project’s next-generation HURD kernel.

If you need guaranteed support, some of Debian’s developers offer support contracts—and so do many third parties. Because we include only Free Software in Debian, you can be assured that you’ll never be “locked in” to any single support vendor—you can always take your support or customisation requirements elsewhere.

Regain your freedom—use free software!

Debian GNU/Linux

Find out more at: http://www.uk.debian.org/

If you’re not interested in this, please give it to someone who might be, since being a non-profit organisation means that we have to try to make the best use of the limited funds available. Thanks.

Debian GNU/Linux technical details:

Current version:  Debian GNU/Linux 2.2 (codenamed “potato”)
Kernel version:  2.2.19 (by default, but you can use any recent kernel)
C Library:  GNU Libc 2.1.3 (will also run glibc 2.0 and old libc5 applications)

Package management: DPKG (Debian Package Management System)

with APT—Debian’s next-generation package-handling tool

The entire system, including the kernel and “init”, is installed via DPKG, so everything can be upgraded very simply. You can upgrade to a new version of Debian without reinstalling from scratch—in fact you don’t even need to shut down the system!

Software included: Over 3950 packages filling three CDs—more than any other distribution. Only the first CD is needed for a fully usable installation. All the standard Linux software is here, and much more. Full source code occupies three optional extra CDs. All software in Debian meets the Debian Free Software Guidelines. Debian is Open Source (the Open Source Definition is based on the Debian Free Software Guidelines)

Debian won’t make your choices for you: we include a variety of components—window managers, email transports, web servers, printing subsystems, etc. You choose which ones meet your needs!

Debian is fully compatible with all other popular Linux distributions. We are actively participating in the Linux Standard Base project, an effort to assure this compatibility for the future. You can install packages from distributions such as Red Hat and Slackware using “alien”, but this is rarely needed, as Debian comes with more prepackaged software than any other distribution. Packages designed specifically for Debian can take advantage of some novel and unique integration strategies, such as the Debian menu system, alternatives system and documentation registry.

Getting Debian source code

Debian is free software—and so that you can modify it to your needs, we distribute full source code. In case you got your Debian binaries without the corresponding source code, you can order a set of three source code CDs from any one of several vendors. For a list, see:

http://www.uk.debian.org/distrib/vendors

Alternatively, you can order from the following suppliers:

Steve McIntyre URL: http://www.chiark.greenend.org.uk/%7Estevem/DebianCD/ Email: stevem@chiark.greenend.org.uk

Philip Hands Phone: 01189 545 656 URL: http://www.hands.com/debiangold.html Email: debiangold@hands.com

Charles Briscoe-Smith Phone: 020 8542 0505 FAX: 020 8287 0637 Email: charles@briscoe-smith.org.uk